OpenStudy (cocoacat):

How would I go about a problem like this?: 2^x = 3^(x-1) I know that usually you try to get the same base but I don't think that's possible here.

OpenStudy (epoweritheta):

bro why your question shows high ??? in red color ??

OpenStudy (cocoacat):

I don't know how to turn it off!

OpenStudy (epoweritheta):

oh...is it taking away you owl-money ??/ okay bring down the \[3^{x}\] over to this side to the denominator

OpenStudy (epoweritheta):

try and tell me

OpenStudy (cocoacat):

I don't have owl money 0-0 (2^x)/(3^(x-1)) = 1

OpenStudy (epoweritheta):

now think how to solve ...

OpenStudy (epoweritheta):

bro keep thinking i will be back in 10 mins ... have to eat :P

OpenStudy (cocoacat):

Ok!

OpenStudy (unklerhaukus):

\[2^x = 3^{x-1}\]Take the log to base 2 on both sides \[\log_2(2^x) = \log_2(3^{x-1})\]Simplify the left hand side \[x = \log_2(3^{x-1})\] Apply the change of base formula \(\boxed{\log_b(a)=\frac{\log_c(a)}{\log_c(b)}}\), (with \(c = 3\) as the new base) \[x =\frac{\log_3(3^{x-1})}{\log_3(2)}\] simplify \[\qquad \vdots\]

OpenStudy (mathmath333):

yup

OpenStudy (epoweritheta):

ooohh.... god... bro multiply and divide by 3 yielding \[3*(2/3)^x=1\] now you can solve it